OpenStudy (anonymous):

PLEASEEEE HELP MEEE!!!! 2 POINTS AWAY FROM AN A!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

I NEED HELP ON THE LAST THREE PROBLEMS

OpenStudy (jhannybean):

Could you possibly use the draw tool or write up your problem? I cannot open your link :(

Nnesha (nnesha):

@Jhannybean :P:DDDDD

Nnesha (nnesha):

2nd page :)

OpenStudy (anonymous):

if you still cannot open the link o will draw out the problem @Jhannybean

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

@triciaal @DanJS

OpenStudy (triciaal):

sorry I don't know this right now. let me solicit some help.

OpenStudy (anonymous):

anything is appreciated @triciaal

OpenStudy (triciaal):

@mathstudent55, @jim_thompson5910, @BPDlkeme234 is anyone available to help?

OpenStudy (anonymous):

i dont think they are online right now. thank you for trying to help me. @triciaal

OpenStudy (triciaal):

f(t) = 1/30(t) + graphs equal at 30 but not at 15 changes by y/30 t = 15 f(15) = 0.5 t = 3*15 f(45) = 1/30(3t) - 0.2 f(5*15) = 1/30(5t) + 0.2 reminds me of sine curve f(t) and g(t) equal on changes of 30 but not on t even multiples has same value

OpenStudy (triciaal):

@Jhannybean can you help please?

OpenStudy (triciaal):

@perl ?

OpenStudy (anonymous):

for those questions about rainfall, 1. yes the graph is invertible because the graph is one to one. It passes both the horizontal and vertical line test.

OpenStudy (anonymous):

2. it asks you to find f(45), or in other words, the cm of rain when t = 45 minutes. If you look at the chart, when t is 45, R is 1.3. So f(45) = 1.3cm

OpenStudy (anonymous):

@jacquelyn1237 Thank you !!!!!!!!!!!!!!!!!!!!!! do you know 3?

OpenStudy (anonymous):

3. To find \[f^{-1} (4.2)\] the f^-1 just means inverse. So now you're looking for when 4.2 cm of rain has fallen, how much time has passed

OpenStudy (anonymous):

So on the chart it'd be 105 minutes :)

OpenStudy (anonymous):

Hope that helped :)

OpenStudy (anonymous):

so the answer for 3 is basiclly 105 minutes? @jacquelyn1237

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you i really appreciate you helping me is there anything i can do to thank you? @jacquelyn1237

OpenStudy (anonymous):

Haha you're welcome, but nah, it's cool you don't have to do anything ;)

OpenStudy (anonymous):

thanks again @jacquelyn1237